# Measuring gravitational lensing time delays with quantum information processing

### v1

@@ VII.1 Example setup
If we consider an example case of using the two Keck telescopes atop Mauna Kea, which have a total collecting area of 152 m2, the resulting signal and background yields are $\approx 66$ and $\approx 105$ per flare, respectively. As a result, we see from Fig. 5 that Keck’s collecting area is insufficient to make a confident mass measurement with one flare alone. As indicated in Fig. 7, for Keck, combining $\gtrsim 10$ flares would be necessary to measure $\Delta t$ with $95\%$ confidence. Given our fiducial flare rate, this corresponds to a roughly 15 day observation, which while long, is not unreasonable. Additionally, it is worth noting that while we have restricted ourselves to studying 1-minute flares that produces $10^{31}$ ergs, the observation of a single 10-minute flare ($n_{\text{sig}}$ = 660) at the same temperature would allow the mass to be estimated at 95% confidence, as can be seen in Fig. 7. As such, the 15-day estimate is likely an overestimate. From Eq. 89, we see that for a 15-day observation, this corresponds to sensitivity lens masses above $\approx 15\,M_{\text{Jup}}$.
@@
The prospects are even better for next-generation extremely large telescopes like the currently under construction Extremely Large Telescope Padovani and Cirasuolo (2023). With a collecting area of 978 m2, the per-flare photon yield would be $\approx 426$ for signal photons and $\approx 677$ for background photons. We see from Fig. 5 that this value means that photons from a single flare are sufficient for us to measure $\Delta t$ with $95\%$ confidence. This would enable the observation of lensing events with duration at least $\approx 1.5$ days, or by Eq. 89, lens masses above $\approx 0.2M_{\text{Jup}}$. Note that much below this mass, the events would become so short that it would be challenging for microlensing surveys to identify that microlensing is occurring in time for us to reorient our telescope. As such, increasing the collecting area of the telescope does not appreciable improve upon the lower mass limit.
Figure 6: Numerical and analytical results for the flare number $m$ necessary to yield a detection as a function the number of signal photons $n_{\mathrm{sig}}$ per flare under various settings. All curves are the analytically-derived sufficient conditions (which are not the optimal limits, see Section VII.2), and points represent the sufficient conditions obtained from numerical simulation, which are closer to optimality. The black curves and points are the baseline setting discussed in our M-dwarf flare observation proposal: $T/t_{c}=10^{4},Q=0.4$, and confidence level is $95\%$. The red curves and points show variations on these three parameters, with one parameter varied in each plot. All simulations are based on magnification $A=1.34$. In (a), we compare the flare number requirement for different confidence levels ($95\%$ and $50\%$). In (b), we alter the signal-to-noise ratio $Q$ from $0.4$ to $0.6$. In (c) we change $T/t_{c}$ from $10^{4}$ to $10^{2}$.
@@
We see that with existing and near-future telescopes, our proposed method could potentially measure the mass of lenses with masses greater than roughly that of Jupiter. The mass range of isolated objects near and above $M_{\text{Jup}}$ is an exceptionally interesting range to explore, as it is currently poorly understood whether the dominant contribution in this mass range arises from free-floating planets or sub-stellar objects Miret-Roig (2023). Above $13\,M_{\text{Jup}}$, the majority of nonluminous lenses would be dim brown dwarfs that are otherwise unobservable, providing a unique way of building a brown dwarf mass function at Galactic distances. Above $\approx 0.1M_{\odot}$, direct mass measurements of observed stellar lenses would allow better calibration of mass-luminosity relations as well as the discovery and characterization of compact objects such as neutron stars and white dwarfs. Finally, at super-solar masses ($M\gtrsim M_{\odot}$), this technique would provide the opportunity to measure the masses of isolated black holes in the mass range probed by LIGO’s observation of black hole mergers Abbott et al. (2023). Note that observations of this duration would also provide a complementary measurement that would help further break lensing degeneracies, even if, for these longer events, orbital parallax can be detected in the light curve. This technique would also enable the direct mass measurement of primordial black holes Carr and Hawking (1974); Carr and Kühnel (2020) and other hypothesized macroscopic dark matter candidates if the dark matter is in fact composed of such objects. As such, if successful, the application of this technique as a follow-up strategy for microlensing surveys would provide the opportunity to do interesting science across many different sub-fields of astronomy. Additionally, as mentioned above, even a non-detection would provide new insight into the temperature and spatial scale of flare emission regions on M dwarfs.
@@
It is worth noting that in the above analysis, we assumed all photons received from the stellar flare in the 1-minute window are in the state close to the $\rho(\Delta t)$ state in Eq. (21), i.e., a superposition of two light paths with stable relative phase and amplitude, even after experiencing dust extinction and passing through the interstellar medium. We justify this robustness assumption in Section VIII.
@@ VII.2 Combination of multiple flares
When the number of photons that can be received per flare is small, we consider using photons from multiple flares that happen in different areas of the star to make a joint analysis. We can safely assume that the durations of the flares do not overlap. Since the size of the M dwarf is generally much greater than the size limit set by the finite-source effect, the difference in $\Delta t$ between different flares is generally much greater than $2\pi/\omega_{0}$. However, we can also assume that this difference is much smaller than $t_{c}$, since $t_{c}$ is a tunable quantity and can be arbitrarily large, in principle.
Let $n$ denote the expected number of photons received per flare and $m$ be the number of flares. For simplicity, suppose that all flares yield exactly $n$ photons. We consider a realistic scenario that takes into account the signal-to-noise ratio $Q$ and magnification $A$. Recall that the score function of $\tau$ for the frequencies $\nu_{i,j}$ (where $j\in\{1,2,\dots,n\}$) corresponding to the $i$th flare is $f(\tau,\nu_{i,1},\dots,\nu_{i,n})=\sum_{j=1}^{n}\cos(\nu_{i,j}\tau)$ with
@@
where $\Delta t_{i}$ is the time delay for the $i$th flare. Since the delays for different flares can vary significantly, the same $\tau$ may correspond to drastically different values of $\cos(\omega_{0}(\Delta t_{i}-\tau))$ for different $i$. We instead introduce a new score function for the $i$th flare, denoted by $f_{\mathrm{c},i}$, and a score function for the combination of all $m$ flares, denoted by $G_{\mathrm{\Delta t_{0}}}$:
$\quad\mathbb{E}[\exp(\mathrm{i}\nu_{i,j}\tau)]$ (94)
$=\int_{-\infty}^{\infty}e^{\mathrm{i}\nu_{i,j}\tau}p_{A}(\nu_{i,j}|\Delta t)d\nu_{i,j}$
$=\sum_{j=1}^{n}\mathopen{}\mathclose{{\left(\mathbb{E}[R_{j}^{2}]+\mathbb{E}[I_{j}^{2}]}}\right)$
$=\sum_{j=1}^{n}\mathbb{E}[R_{j}^{2}+I_{j}^{2}]d=n.$
$\mathbb{E}[f_{\mathrm{c},i}]$ $=\mathopen{}\mathclose{{\left|\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{j=1}^{n}\exp(\mathrm{i}\nu_{i,j}\tau)}}\right]}}\right|^{2}+$ (96)
$\quad\operatorname{Var}\mathopen{}\mathclose{{\left(\sum_{j=1}^{n}\exp(\mathrm{i}\nu_{i,j}\tau)}}\right)$
$=\frac{n^{2}Q^{2}\gamma_{A}^{2}}{4}+\sum_{j=1}^{n}\mathbb{E}[R_{j}^{2}+I_{j}^{2}]$
$\quad-\sum_{j=1}^{n}\mathopen{}\mathclose{{\left[\mathbb{E}[R_{j}]^{2}-\mathbb{E}[I_{j}]^{2}}}\right]$
$=\frac{n^{2}Q^{2}\gamma_{A}^{2}}{4}-\frac{nQ\gamma_{A}^{2}}{4}+n$
$=\frac{\gamma_{A}^{2}}{4}(n_{\mathrm{sig}}^{2}-n_{\mathrm{sig}})+n$
$\quad\mathbb{E}[f_{\mathrm{c},i}(\tau)]$ (97)
@@
$\quad\mathbb{E}[G_{\mathrm{\Delta t_{0}}}(\tau)]$ (98)
[…]
$\quad\Pr[f_{\mathrm{c},i}(\tau)-n>\epsilon]$ (99)
$\leq\exp(-\frac{(\sqrt{\epsilon+n}-\sqrt{n})^{2}}{2n})$
$=\exp(-\frac{\epsilon}{2n}-1+\sqrt{1+\frac{\epsilon}{n}}).$
The above upper bound converges to $\exp(-\frac{\epsilon^{2}}{8n^{2}})$ when $\epsilon\ll n$ and becomes $\exp(-\frac{\epsilon}{2n})$ when $\epsilon\gg n$. This implies that $f_{\mathrm{c},i}$ is approximately a sub-exponential random variable with parameters $(4n^{2},n)$. Therefore, as a sum of sub-exponential random variables, Bernstein’s inequality Chafaï et al. (2012) gives the following bounds for $G_{\mathrm{\Delta t_{0}}}$:
$\quad\Pr[G_{\mathrm{\Delta t_{0}}}(\tau)>G_{\mathrm{th}}]$ (100)
[…]
and for $n_{\mathrm{sig}}$ close to $2$, the sufficient condition is
Again, we emphasize that the above derivation only gives the sufficient condition for a successful time delay measurement, or an upper bound (in fact, a loose upper bound) of $m$ for any given $(n_{\mathrm{sig}},T/t_{c},Q,A)$ parameters. The actual lower bound for $m$ will likely have different constant factors, but will never be higher than the upper bounds given by the sufficient-condition analysis. Indeed, we perform numerical simulations to find the number of flares actually needed for various realistic parameters, and the results are significantly less than the theoretical upper bounds. We present the numerical results and their comparison with theoretical sufficient conditions in Fig. 6. We also give one realistic example (which is the potentially most interesting example) for the fiducial M-dwarf flare setting considered in Section VII.1 where $A=1.34$ (hence $\gamma_{A}=0.666$) and $Q=0.386$. If we set $T/t_{c}=10000$, use the Extremely Large Telescope (which gives $n_{\mathrm{sig}}=426$), and aim for $95\%$ confidence, then the sufficient condition given by our the analytic bound is $m\geq 3$, while numerical simulation shows that even $m=1$ works.
To give an intuitive demonstration of how our algorithm works, in Fig. 7, we present the score function values from the numerical simulation of one successful time-delay measurement. The results are from (the simulation of) the same fiducial observation of M dwarf flares but using Keck rather than the Extremely Large Telescope. It combines 5 flares with 66 signal photons and 103 noise photons per flare. The fact that $\Delta t$ is successfully found with only 5 flares in this simulation is not surprising, because according to Fig. 6(a), 64 photons with $Q=0.4$ need only $6$ flares to achieve $50\%$ success probability.
@@ VIII Robustness of photonic states against the medium
As mentioned in the previous section, the interstellar medium may cause various effects on the observation, including photon loss due to dust extinction Saha et al. (2019) (corresponding to the imaginary part of the refractive index) and additional phases imprinted by gases (corresponding to the real part of the refractive index). In this section, we perform a thorough analysis of the robustness of our observation scheme (or, more fundamentally, the $\Delta t$-information in the optical signal) to both effects. We also briefly discuss the final section of the medium between the source and the telescope, which is our atmosphere, in Section VIII.3, and claim that it does not affect the stability of our measurement.
@@ VIII.1 Dust extinction
The increased telescope size requirement due to the photon loss caused by dust extinction has been taken into account in the analysis of the feasibility of our example setup in Eqs. (90, 91) in Section VII.1. However, since our delay-finding approach is based on the interference of two branches of the same particle, we additionally require each photon received from the source to be in one of the pure states described by Eq. (20) (or a state close to it) with a random $t_{0}$. This means that the state must be a superposition of two branches with comparable and stable weight with a stable relative phase between them. In this subsection, we discuss the potential effect of dust on the weights in the superposition.
@@
Fortunately, our close examination shows that the reality is much closer to the beam splitter model as long as the size of dust particles are sufficiently small compared to the radius of the telescope. For meter-level telescope, this is almost certainly true because Ref. Mathis et al. (1977) suggests that $1\,\mathrm{\mu m}$ is a reasonable estimate of particle size. We provide a detailed analysis in Appendix A, in which we model the microscopic dust particle configuration as a binary tree coloring problem and derive an explicit relation between the level of decoherence and the size of dust particles.
@@ VIII.2 Variations in refractive index
and scales linearly with with $D_{S}$, the distance to the source Armstrong et al. (1995); You et al. (2007). The structure function also allows for the estimation of a root-mean-square variability of $\mathrm{DM}/dt$ for a given scale as $\sigma_{\text{DM}}\approx\sqrt{D_{\text{DM}}(\tau)}/\tau$ Donner et al. (2020).
@@
Though there are few distant pulsars with well-measured structure functions, local pulsars Donner et al. (2020) indicate typical values of $D_{\text{DM}}(1000\,\text{day})\approx 1\times 10^{-6}\penalty 10000\ \frac{\text{pc}^{2}}{\text{cm}^{6}}$ and $D_{S}\approx 1$ kpc. Using this, and evaluating Eq. (108) for typical values of interest for our proposal, we find
@@
over 1 minute, hence is negligible for the timescales of interest to our proposal. Even taking the pulsar with the highest $D_{\mathrm{DM}}(\mathrm{1000\,day})=2.2\times 10^{-4}\,\mathrm{pc}^{2}\,\mathrm{cm}^{-6}$ value from Donner et al. (2020), we still get only $\sigma_{\phi}=9.9\times 10^{-7}$ (average value for this dataset is $\sigma_{\phi}=0.98\times 10^{-7}$).

### v2

@@ VII.1 Example setup
If we consider an example case of using the two Keck telescopes atop Mauna Kea, which have a total collecting area of 152 m2, the resulting signal and background yields are $\approx 132$ ($\approx 198$) and $\approx 210$ ($\approx 315$) per flare, respectively, assuming each flare has $2$-minute (33-minute) duration. As a result, we see from Table 2 that Keck’s collecting area is insufficient to make a confident mass measurement with one flare alone. As indicated in Table 2, for Keck, we would need to combine $\gtrsim 7$ flares with $3$-minute duration to measure $\Delta t$ with $70\,\%$ confidence when $T/t_{c}=10^{8}$, corresponding to $M\approx 10\,M_{\mathrm{Jup}}$. Given our fiducial flare rate, this corresponds to a roughly 11-day observation, which is not unreasonable since the typical duration of a lensing event with $10\,M_{\mathrm{Jup}}$ lens mass is $\gtrsim 13\,\mathrm{days}$, according to Eq. (92). Additionally, it is worth noting that while we have restricted ourselves to studying 2-minute or 3-minute flares that produce $10^{31}$ ergs, the observation of a single 10-minute flare ($n_{\text{sig}}$ = 660) at the same temperature would allow the mass to be estimated with very high confidence for various settings of $T/t_{c}$ and $Q$, as suggested by Fig. 5.
@@
The prospects are even better for next-generation extremely large telescopes like the currently under construction Extremely Large Telescope [90]. With a collecting area of 978 m2, the per-flare photon yield for a 1-minute (2-minute) flare would be $\approx 426$ ($\approx 852$) for signal photons and $\approx 677$ ($\approx 1352$) for background photons. We see from Fig. 5 that in this scenario, photons from a single $1$-minute flare are sufficient to measure $\Delta t$ with $70\,\%$ confidence for $T/t_{c}=10^{6}$, and a single 2-minute flare would enable $\Delta t$-measurement for $T/t_{c}>10^{8}$ with $95\%$ confidence. This would enable the observation of lensing events with duration at least $\approx 1.5$ days, or by Eq. (92), lens masses above $\approx 0.2M_{\text{Jup}}$. Note that much below this mass, the events would become so short that it would be challenging for microlensing surveys to identify that microlensing is occurring in time for us to reorient our telescope. As such, increasing the collecting area of the telescope does not appreciably improve upon the lower mass limit.
@@
We see that with existing and near-future telescopes, our proposed method could potentially measure the mass of lenses with masses greater than roughly that of Jupiter. The mass range of isolated objects near and above $M_{\text{Jup}}$ is an exceptionally interesting range to explore, as it is currently poorly understood whether the dominant contribution in this mass range arises from free-floating planets or sub-stellar objects [12]. Above $13\,M_{\text{Jup}}$, the majority of nonluminous lenses would be dim brown dwarfs that are otherwise unobservable, providing a unique way of building a brown dwarf mass function at Galactic distances. Above $\approx 0.1M_{\odot}$, direct mass measurements of observed stellar lenses would allow better calibration of mass-luminosity relations as well as the discovery and characterization of compact objects such as neutron stars and white dwarfs. Finally, at super-solar masses ($M\gtrsim M_{\odot}$), this technique would provide the opportunity to measure the masses of isolated black holes in the mass range probed by LIGO’s observation of black hole mergers [61]. Note that observations of this duration would also provide a complementary measurement that would help further break lensing degeneracies, even if, for these longer events, orbital parallax can be detected in the light curve. This technique would also enable the direct mass measurement of primordial black holes [91, 92] and other hypothesized macroscopic dark matter candidates if the dark matter is in fact composed of such objects. As such, if successful, the application of this technique as a follow-up strategy for microlensing surveys would provide the opportunity to do interesting science across many different sub-fields of astronomy. Additionally, as mentioned above, even a non-detection would provide new insight into the temperature and spatial scale of flare emission regions on M dwarfs.
@@
It is worth noting that in the above analysis, we assumed all photons received from the stellar flare in the 1-minute or 2-minute window are in the state close to the $\rho(\Delta t)$ state in Eq. (21), i.e., a superposition of two light paths with stable relative phase and amplitude, even after experiencing dust extinction and passing through the interstellar medium. We justify this robustness assumption in Section VIII.
@@ VII.2 Combination of multiple flares
When the number of photons that can be received per flare is smaller, we consider using photons from multiple flares that happen in different areas of the star to make a joint analysis. We can safely assume that the durations of the flares do not overlap. Since the size of the M dwarf is generally much greater than the size limit set by the finite-source effect, the difference in $\Delta t$ between different flares is generally much greater than $2\pi/\omega_{0}$. Since the photons result from broadband emission, this difference is also much greater than the coherence time $t_{c}$. Note that if this difference is smaller than the coherence time, this will be the narrow-band scenario, in which combining multiple flares may have better performance. Although the narrow-band case is less realistic, we present its strategy in Appendix B in case this is of independent interest.
Let $n$ denote the expected number of photons received per flare and $m$ be the number of flares. For simplicity, suppose that all flares yield exactly $n$ photons. We consider a realistic scenario that takes into account the signal-to-background ratio $Q$ and magnification $A$. Recall that the score function of $\tau$ for the frequencies $\nu_{i,j}$ (where $j\in\{1,2,\dots,n\}$) corresponding to the $i$th flare is $f(\tau,\nu_{i,1},\dots,\nu_{i,n})=\sum_{j=1}^{n}\cos(\nu_{i,j}\tau)$ with
@@
where $\Delta t_{i}$ is the time delay for the $i$th flare.
Since the delays for different flares may vary significantly, we aim to find the smallest possible time window of $\tau$ that almost all $\Delta t_{i}$s fall into. An intuitive way to determine the size of the time window is to match the standard deviation of the $\Delta t_{i}$s, denoted by $\delta_{\Delta t,\mathrm{f}}$. This quantity is determined by the linear size of the host M dwarf, which is typically between $0.1\,R_{\odot}$ and $0.7\,R_{\odot}$ (note that in the fiducial case in Table 1, we use $0.2\,R_{\odot}$). This corresponds to approximately $\delta_{\Delta t,\mathrm{f}}\in[400t_{c},20000t_{c}]$. If we take $\delta_{\Delta t,\mathrm{f}}=1000t_{c}$ as a fiducial setting and assume we have prior knowledge that $T=10^{-8}\,\mathrm{s}$ (corresponding to $T/t_{c}=10^{6}$), then our task is to find the correct window containing almost all $\Delta t_{i}$s among the $T/\delta_{\Delta t,\mathrm{f}}=1000$ time windows.
To design an algorithm for the above task, we propose a score function for the combination of multiple flares. A straightforward idea is directly summing the absolute values of every flare’s score function, and checking which time window has the highest mean score or maximum score. Note that such a sum of score functions is unlikely to create a significantly large value at one specific $\tau$ when the number of flares is small, because every peak in the score function only has width $t_{c}$ and the probability that two peaks overlap is tiny.
We generalize the above idea to give the following sum-$L_{p}$ score functions. Letting $p\in[0,\infty]$ be the order of the $L_{p}$ norm, and $j\in\{0,1,\dots,T/\delta_{\Delta t,\mathrm{f}}-1\}$ be the index of time window, we find that
$G_{L_{p}}(j)\coloneqq\sum_{i=0}^{m-1}\mathopen{}\mathclose{{\left[\sum_{k=0}^{\delta_{\Delta t,f}/t_{c}-1}\mathopen{}\mathclose{{\left|f_{i}(j\delta_{\Delta t,\mathrm{f}}+kt_{c})}}\right|^{p}}}\right]^{1/p}$ (96)
is our score function for the $j$th window (corresponding to $\tau\in[j\delta_{\Delta t,\mathrm{f}},(j+1)\delta_{\Delta t,\mathrm{f}}]$). The time window inferred from $m$ flares using the above score function is
$j_{\mathrm{opt}}=\arg\max_{j}G_{L_{p}}(j).$ (97)
Note that $p$ can be freely chosen to adjust our focus—whether we emphasize the average score of the window or its peak score. For $p=1$, this is the standard summation of score function values, while for $p=\infty$, we take the maximum among all score function values inside the window. To choose the best $L_{p}$ norm, we perform a numerical analysis for a small-scale example where $m=10$ flares are combined, each flare has $n_{\mathrm{sig}}=132$, with $A=1.34$, $Q=0.4$, $T/t_{c}=10^{4}$, and $\delta_{\Delta t,\mathrm{f}}/t_{c}=400$. We summarize the results in Table 3.
$p$ confidence level
11 $52\pm 7\,\%$
22 $64\pm 7\,\%$
$\infty$ $96\pm 3\,\%$
Table 3: Results for numerical simulation of the performance of sum-$L_{p}$ score functions for $n_{\mathrm{sig}}=132$, $m=10$, $T/t_{c}=10^{4}$, and $\delta_{\Delta t,\mathrm{f}}/t_{c}=400$. The confidence level is the probability for the algorithm to find the correct time window containing the $\Delta t_{i}$s.
Clearly, $G_{L_{\infty}}$ is the score function with best performance. Intuitively, if $\tau$ is close to $\Delta t_{i}$, $f_{i}(\tau)$ will be a peak with high probability, hence the window containing correct $\Delta t_{i}$s will more likely have a higher peak value.
@@ 22 $64\pm 7\,\%$
Indeed, our numerical simulation (see Table 2) shows that our flare combination strategy offers significant improvement over the single-flare case even if combining just a limited number of flares. These numerical results are based on more realistic settings where $T/t_{c}\in\{10^{5},10^{6},10^{7},5\times 10^{7}\}$ and $\delta_{\Delta t,\mathrm{f}}/t_{c}=10^{3}$. Note that each $T/t_{c}$ corresponds to a typical duration of the lensing event, $t_{E}$, and subsequently a typical number of flares to expect, $m_{\mathrm{typ}}$. When the number of flares needed is smaller than $m_{\mathrm{typ}}$ for a certain confidence level, then with high probability one can get sufficiently many flares to obtain a $\Delta t$ estimation with that confidence. Observe that for a certain $n_{\mathrm{sig}}$, the required $m$ increases slowly while $T/t_{c}$ increases exponentially, which aligns with the result that the total number of photons ($n_{\mathrm{sig}}m/Q$) is proportional to $\log(T/t_{c})$. Therefore, typically, our flare combination strategy works better with longer time delay. Indeed, Keck is capable of measuring $\Delta t$ with $98\,\%$ confidence when $T/t_{c}=5\times 10^{7}$.
We emphasize that the numerical results presented in this subsection are based on a specific fiducial setting. In practice, there are more realistic factors affecting the algorithm’s performance. For instance, the time delay will drift during the whole lensing event due to the transverse velocity. However, we claim that this effect, along with many other possible issues, can be taken into account in algorithm design. For the drift of time delay, one can use the magnification data and Fig. 2 to model the change of $\Delta t$ and fit the frequency measurement outcomes to recover the entire time-delay curve. A more detailed exploration of possible variants of this algorithm is beyond the scope of this paper.
@@ VIII Robustness of photonic states against the medium
As mentioned in the previous section, the interstellar medium may cause various effects on the observation, including photon loss due to dust extinction [87] (corresponding to the imaginary part of the refractive index) and additional phases imprinted by gases (corresponding to the real part of the refractive index). In this section, we perform a thorough analysis of the robustness of our observation scheme (or, more fundamentally, the $\Delta t$-information in the optical signal) to both effects. We also briefly discuss the final section of the medium between the source and the telescope, which is our atmosphere, in Section VIII.3, and claim that it does not affect the stability of our measurement.
@@ VIII.1 Dust extinction
The increased telescope size requirement due to the photon loss caused by dust extinction has been taken into account in the analysis of the feasibility of our example setup in Eqs. (93, 94) in Section VII.1. However, since our delay-finding approach is based on the interference of two branches of the same particle, we additionally require each photon received from the source to be in one of the pure states described by Eq. (20) (or a state close to it) with a random $t_{0}$. This means that the state must be a superposition of two branches with comparable and stable weight with a stable relative phase between them. In this subsection, we discuss the potential effect of dust on the weights in the superposition.
@@
[…]
Fortunately, our close examination shows that the reality is much closer to the beam splitter model as long as the size of dust particles are sufficiently small compared to the radius of the telescope. For meter-level telescope, this is almost certainly true because Ref. [93] suggests that $1\,\mathrm{\mu m}$ is a reasonable estimate of particle size. We provide a detailed analysis in Appendix A, in which we model the microscopic dust particle configuration as a binary tree coloring problem and derive an explicit relation between the level of decoherence and the size of dust particles.
@@ VIII.2 Variations in refractive index
and scales linearly with with $D_{S}$, the distance to the source [94, 95]. The structure function also allows for the estimation of a root-mean-square variability of $\mathrm{DM}/dt$ for a given scale as $\sigma_{\text{DM}}\approx\sqrt{D_{\text{DM}}(\tau)}/\tau$ [96].
@@
Though there are few distant pulsars with well-measured structure functions, local pulsars [96] indicate typical values of $D_{\text{DM}}(1000\,\text{day})\approx 1\times 10^{-6}\penalty 10000\ \frac{\text{pc}^{2}}{\text{cm}^{6}}$ and $D_{S}\approx 1$ kpc. Using this, and evaluating Eq. (102) for typical values of interest for our proposal, we find
@@
over 1 minute, hence is negligible for the timescales of interest to our proposal. Even taking the pulsar with the highest $D_{\mathrm{DM}}(\mathrm{1000\,day})=2.2\times 10^{-4}\,\mathrm{pc}^{2}\,\mathrm{cm}^{-6}$ value from [96], we still get only $\sigma_{\phi}=9.9\times 10^{-7}$ (average value for this dataset is $\sigma_{\phi}=0.98\times 10^{-7}$).